OpenStudy (anonymous):

Louisa has five shirts (white, gray, black, green, and blue), four pairs of pants (black, khaki, navy, and jeans), and two pairs of shoes (flip flops and sneakers). How many different outfits can she create?

OpenStudy (anonymous):

5! x 4! x 2!

OpenStudy (anonymous):

40

OpenStudy (anonymous):

no. factorial.! 5! = 5*4*3*2*1 4!=4*3*2*1

OpenStudy (anonymous):

how does that work?

OpenStudy (anonymous):

the no. of different outfit would be 5! x 4! x2! =(5*4*3*2*1)x(4*3*2*1)x(2*1) =(120)(24)(2) =5760

OpenStudy (anonymous):

okay....